An entrepreneur bought computers for his business, twice as many laptops as fixed equipment. The fixed equipment is worth $14,500.00 and the laptop $5,800.00. If he spent a total of $130,500.00, what was the total amount of equipment he purchased?
Answer by math_tutor2020(3817)


x = number of fixed equipment
2x = number of laptops, based on the last part of the first sentence

The second sentence "The fixed equipment is worth $14,500.00 and the laptop $5,800.00" means
5800*2x = 11600x = cost of all the laptops
14500x = cost of all the fixed equipment

Those subtotals must add to $130,500.00 which is the total spent.

11600x+14500x = 130500
26100x = 130500
x = 130500/26100
x = 5
The person bought 5 items of fixed equipment
He also bought 2x = 2*5 = 10 laptops.

In total, he bought 5+10 = 15 items.